# Confabulation dynamics in a reservoir computer: Filling in the gaps with untrained attractors

### v2

@@ II.2 Training RC to reconstruct an attractor
with respect to $\textbf{W}_{out}$, where $||\cdot||_{2}$ is the Euclidean norm, $l^{*}=t_{\text{listen}}/\tau$ and $t^{*}=t_{\text{train}}/\tau$, $\bm{r}[i]$ and $\bm{u}[i]$ are discrete-time samples of the continuous-time variables $\bm{r}(t)$ and $\bm{u}(t)$ at discrete time $i=t/\tau$. The corresponding time series, $\left[\bm{r}[i]\right]_{i=l^{*}}^{t^{*}}$ and $\left[\bm{u}[i]\right]_{i=l^{*}}^{t^{*}}$, are constructed by sampling $\bm{r}(t)$ and $\bm{u}(t)$ at intervals of length $\tau$ (discrete-time intervals of length $1$). $\beta$ is the regularisation parameter and the purpose of the $\beta\,||\textbf{W}_{out}||_{2}^{2}$ term in Eq. (8) is to modify the linear least-squares regression to reduce the magnitudes of elements in $\textbf{W}_{out}$ in order to discourage overfitting.
Minimising Eq. (8) involves computing its partial derivative with respect to $\textbf{W}_{out}$ and setting the resulting expression equal to 00. The $\textbf{W}_{out}$ that minimises Eq. (8) is given by
@@
$\dot{\hat{\bm{r}}}(t)$ $=\gamma\left[-\hat{\bm{r}}(t)+\tanh{\left(\textbf{M}\,\hat{\bm{r}}(t)+\sigma\textbf{W}_{in}\textbf{W}_{out}\,\bm{q}(\hat{\bm{r}}(t))\right)}\right],$ (14)
$\hat{\bm{r}}\left(0\right)$ $=\bm{r}\left(t_{train}\right),$ (15)
@@ II.3 Training a RC to reconstruct attractors at different parameter settings
$\dot{\hat{\bm{r}}}(t)=\gamma\left[-\hat{\bm{r}}(t)+\tanh{\left(\textbf{M}\hat{\bm{r}}(t)+\sigma\textbf{W}_{in}\textbf{W}_{out}^{C}\,\bm{q}(\hat{\bm{r}}(t))+\bm{b}\right)}\right].$ (20)
@@ II.4 Defining untrained and generated attractors
Definition 1:An attractor, $\pazocal{Z}$, is considered to be an ‘untrained attractor’ if (i) $\pazocal{Z}$ was not present during training, and (ii) $\pazocal{Z}$ does not originate from a continuous deformation/displacement or bifurcation of a given reconstructed attractor (by varying certain parameters of the given reservoir computer).
@@
Definition 2:An attractor, $\pazocal{G}$, is considered to be a ‘generated attractor’ if (i) $\pazocal{G}$ was not present during training, and (ii) $\pazocal{G}$ originates from a continuous deformation/displacement or bifurcation of a given reconstructed attractor (by varying certain parameters of the given reservoir computer).
@@ II.5.1 Task (i) description
$\begin{array}[]{ccl}\dot{x}_{1}&=&10\left(x_{2}+x_{1}\right),\\ \vspace{-0.3cm}\hfil&&\\ \dot{x}_{2}&=&x_{1}\left(28-x_{3}\right)-x_{2},\\ \vspace{-0.3cm}\hfil&&\\ \dot{x}_{3}&=&x_{1}x_{2}-\frac{8}{3}x_{3},\end{array}$ (26)
@@
To assess how well the closed-loop RC reconstructs $\pazocal{L}$, we use the method outlined in Appendix B. From this we classify the output from a given initial condition as a ‘good’ reconstruction of $\pazocal{L}$, denoted by $\hat{\pazocal{L}}$, a ‘poor’ reconstruction of $\pazocal{L}$, denoted by $\tilde{\pazocal{L}}$. If there is no reconstruction of $\pazocal{L}$, i.e. there is no stable attractor located nearby $\pazocal{L}$ with properties resembling $\pazocal{L}$, we then say that the state of the closed-loop RC approaches an UA, corresponding to complete confabulation. We find a wide variation of dynamics corresponding to $\tilde{\pazocal{L}}$ and later show examples of these attractors in Sec. III and Appendix C.
@@ II.5.2 Task (ii) description
$\begin{array}[]{ccl}\dot{x}_{1}&=&x_{2}\left(1+x_{3}\right),\\ \vspace{-0.3cm}\hfil&&\\ \dot{x}_{2}&=&x_{3}\left(x_{2}-ax_{1}\right),\\ \vspace{-0.3cm}\hfil&&\\ \dot{x}_{3}&=&0.55~{}x_{3}^{2}-x_{2}^{2}.\end{array}$ (32)
@@
In Fig. 2 we show that for decreasing $a$, the period-1 limit cycle undergoes a sequence of period-doubling bifurcations. Therefore, if the parameter-aware closed-loop RC successfully interpolates the dynamics of Eq. (32) between the period-1 and period-4 limit cycles, then the RC should generate a period-2 limit cycle at some intermediate parameter setting between $b$ and $-b$. In addition, the RC should generate a continuous transition between this period-2 limit cycle and the reconstructed period-1 and period-4 limit cycles, where each transition should occur due to a period-doubling bifurcation. Furthermore, the RC should continue to undergo successive period-doubling bifurcations beyond the period-4 limit cycle to produce chaotic dynamics. However, as our results show, it is not so straightforward for the RC to fill the gaps in its knowledge of Eq. (32) by producing GAs that mimic the dynamics shown in Fig. 2.
@@ II.5.3 Task (iii) description
$\begin{array}[]{ccl}\dot{x}_{1}&=&-1.3x_{1}-4\left(x_{2}+x_{3}\right)-x_{2}^{2},\\ \vspace{-0.3cm}\hfil&&\\ \dot{x}_{2}&=&-1.3x_{2}-4\left(x_{3}+x_{1}\right)-x_{3}^{2},\\ \vspace{-0.3cm}\hfil&&\\ \dot{x}_{3}&=&-1.3x_{3}-4\left(x_{1}+x_{2}\right)-x_{1}^{2},\end{array}$ (38)
@@ II.6 Summary of related work
While the presence of UAs in RC systems are being increasingly studied and acknowledged [36, 22, 37, 38, 39], it is quite likely that UAs have lurked in the background of many researchers dealings with RCs but have simply gone unnoticed. For instance, in one of the most influential papers on reservoir computing in recent times, Lu et al. [23], the authors present an interesting result in their Fig. 5 where they show that for a relatively small change in the input scaling parameter, $\sigma$, there is an abrupt transition in the RC’s dynamics from a period-2 limit cycle to the chaotic Lorenz attractor. Based on the results presented in the present paper, it is not too far-fetched to speculate that the limit cycle is an UA that coexists with the reconstructed Lorenz attractor. Other examples of UAs potentially hiding in plain sight can be seen in de Jong et al. [40].
@@ III.1 Frequency of scenarios in Table 1 across 50 different M
Fig. 4 shows that for $\rho=0$, the only attractor present in each $\mathbb{P}$ is an UA. By increasing $\rho$ to $0.05$, there is a dramatic change in $\mathbb{P}$ as the case of a poor reconstruction of $\pazocal{L}$ coexisting with UAs becomes the dominant scenario. For $\rho=0.1$ we see the first occurrences of RCs that are able to reconstruct $\pazocal{L}$, albeit poorly, without any UAs in $\mathbb{P}$. This trend continues for increasing $\rho$ and when $\rho=0.2$ we see the first instances of RCs that are able to provide a good reconstruction of $\pazocal{L}$ without any UAs present in $\mathbb{P}$. By increasing $\rho$ further we see this becomes the dominant scenario for $0.3\lesssim\rho\lesssim 0.7$. Within this range of $\rho$ values, UAs are found for a decreasing number of M. For $0.7\lesssim\rho\lesssim 1.1$, Fig. 4 shows that there is a relatively equal likelihood that the RC will provide either a good or bad reconstruction of $\pazocal{L}$, and that it is much less likely for these reconstructions to coexist with UAs. However, there is a steady increase in the number of RCs that fail to provide any reconstruction of $\pazocal{L}$ within this range of $\rho$ values, thus, resulting in an increase in the UA only scenario. These trends continue for $\rho>1.1$ with poor reconstructions of $\pazocal{L}$ becoming the dominant scenario.
@@
In Fig. 7 we re-present the results of Fig. 4 from the point of view of each M in order to illustrate the changes that occur in each of the corresponding $\mathbb{P}$ when changing $\rho$. We label each M by $\textbf{M}_{i}$, from $i=1$ to $50$, in order of the number of $\rho$ values that the corresponding closed-loop RC in Eq. (14) provides a good reconstruction of $\pazocal{L}$, with $i=1$ being the most amount of $\rho$ values. Fig. 7 provides us with additional information, such as: some random realisations of M quickly overcome the influence of UAs, like $\textbf{M}_{1}$ and $\textbf{M}_{4}$, whereas others require relatively large $\rho$, like $\textbf{M}_{43}$ needing $\rho>0.9$.
@@ III.2 UAs filling in the gaps when reconstruction of $\pazocal{L}$ fails
Fig. 7 shows that when $\rho=0.05$ the majority of the different M matrices are able to provide a poor reconstruction of $\pazocal{L}$, denoted by $\tilde{\pazocal{L}}$, which coexists with at least one UA. In this subsection, we illustrate the most common way in which $\tilde{\pazocal{L}}$ comes into existence and another less common but equally interesting way. In Appendix C we provide additional examples of interesting ways in which the RC reconstructs $\pazocal{L}$.
III.2.1 Lorenz chaotic transient becoming stable
In Fig. 8 we show the most common way that the RC provides a reconstruction of $\pazocal{L}$ when increasing $\rho$ from 00 for one of the matrices $\textbf{M}_{i}$ from our ensemble. In each panel of Fig. 8, the black dotted curves correspond to the output of the closed-loop RC (Eq. (14)) when initialised with the last point of the training, $\bm{r}(t_{train})$. The same colour scheme from Fig. 5 is used to illustrate different UAs that we find. In panel (a) we show that while the closed-loop RC is unable to provide any reconstruction of $\pazocal{L}$, the brief transient behaviour of the RC’s state approaching UA1 resembles Lorenz-like behaviour. In panels (b)–(e) we see that before approaching UA1 or UA2, the duration of this transient activity increases and the transient also grows in its resemblance to $\pazocal{L}$. In panel (f) we no longer see this transient behaviour as the RC is able to provide a poor reconstruction of $\pazocal{L}$. However, this reconstruction has some noticeable features on the right-hand wing ($x_{1}>0$ portion of $\mathbb{P}$) that are not consistent with $\pazocal{L}$ which is why our algorithm classifies this attractor as a poor reconstruction of $\pazocal{L}$. In panels (g)–(i) we see that as $\rho$ increases, UA2 shrinks in size. Interestingly, in panel (j), after UA2 becomes unstable, this activity on the right-hand wing of the reconstructed $\pazocal{L}$ is no longer visible, and this attractor is classified as a good reconstruction of $\pazocal{L}$. The RC maintains its good reconstruction of $\pazocal{L}$ for larger values of $\rho$. Without too much speculation, one reason for this could be that since UA2 is located near $\pazocal{L}$ in this projection, it may also be in $\mathbb{S}$, and therefore the saddle which separates these states may play a role in preventing the RC from achieving a good reconstruction of $\pazocal{L}$, and after UA2 becomes unstable this enables the RC to improve its reconstruction.
In summary, Fig. 8 shows a sequence of events where we see a transition from an unstable reconstruction of $\pazocal{L}$ to a stable reconstruction of $\pazocal{L}$ with a stable UA filling in the gaps. From further investigation we find transitions from a stable UA and unstable reconstruction of $\pazocal{L}$ to an unstable UA and stable reconstruction of $\pazocal{L}$. We also find switching dynamics via transitions from a stable UA and unstable reconstruction of $\pazocal{L}$ to an unstable UA and unstable reconstruction of $\pazocal{L}$ before another transition to an unstable UA and stable reconstruction of $\pazocal{L}$. As we show in Flynn and Amann [28], when two attractors become unstable in quick succession of one another and there are no attractors present in $\mathbb{P}$, then the state of the RC will switch between transients of both attractors in finite time and remain in either transient for different durations of time.
III.2.2 Reconstruction of $\pazocal{L}$ via period-doubling route to chaos
In Fig. 9, we show that by increasing the parameter $\rho$ of $\textbf{M}_{28}$ from 00, the first time the RC reconstructs a Lorenz-like attractor, it reconstructs $\pazocal{L}$ as a limit cycle. In the inset panel we see that this limit cycle undergoes a period-doubling route to chaos before becoming unstable. To provide further insight, in Fig. 10 we illustrate the dynamics of this limit cycle as it undergoes these bifurcations at different $\rho$ values. As indicated by the arrow in Fig. 10 (e), after this reconstruction becomes unstable the state of the RC then approaches UA1 as this is the only stable state in $\mathbb{P}$ for these $\rho$ values. Prior to becoming unstable, this reconstruction grows in its resemblance to $\pazocal{L}$, which is why our algorithm classifies this as a poor reconstruction of $\pazocal{L}$ rather than an UA. At larger $\rho$ values, we see this poor reconstruction of $\pazocal{L}$ reappear and disappear within different ranges of $\rho$ values for $\rho<0.1$ in a similar way to the events shown in Fig. 8. Within this range of $\rho$ values we also find another UA, a fixed point that we denote by UA5, which the state of the RC approaches when reconstruction fails for $0.025\lesssim\rho\lesssim 0.03$. For $0.3250.1<\rho<$ 0.325 we find that the reconstruction exists as a transient and, depending on the value of $\rho$, this transient activity can last for significantly long durations of time. In order to determine whether the RC provides a stable reconstruction of $\pazocal{L}$ within this large range of $\rho$ values, we had to increase $t_{predict}$ to $30,000$. The RC achieves a good reconstruction of $\pazocal{L}$ for $\rho=0.325$ and maintains this for much larger $\rho$ values.

### v3

@@ I Introduction
When discussing confabulation mechanisms in the wider literature, researchers from both the psychology and machine learning communities argue that confabulations may arise when the brain or some learning system is ‘filling in the gaps’ in its memory in cases where an accurate solution is not found for a particular query. For instance, in the APA Dictionary of Psychology, confabulations are defined as “the falsification of memory in which gaps in recall are filled by fabrications that the individual accepts as fact” [26]. In a machine learning context, Nejjar et al. [27] describe confabulations in terms of “…where an LLM [large language model] seemingly fills gaps in the information contained in the model with plausibly-sounding words”. However, to clearly identify and visualise such ‘gaps’ being filled on a network level is challenging due to the complexity of these network systems. In contrast, the value of our RC approach is that it is mathematically tractable and we are able to produce diagrams of intended, learned solutions with gaps that are naturally filled by alternative, sometimes unexpected, solutions. More specifically, our results illustrate that what we observe as confabulations is literally the presence of unexpected solutions (UAs) “filling in the gaps”.
@@ II.2 Training RC to reconstruct an attractor
with respect to $\textbf{W}_{out}$, where $||\cdot||_{2}$ is the Euclidean norm, $l^{*}=t_{\text{listen}}/\tau$ and $t^{*}=t_{\text{train}}/\tau$, $\bm{r}[i]$ and $\bm{u}[i]$ are discrete-time samples of the continuous-time variables $\bm{r}(t)$ and $\bm{u}(t)$ at discrete time $i=t/\tau$. The corresponding time series, $\left[\bm{r}[i]\right]_{i=l^{*}}^{t^{*}}$ and $\left[\bm{u}[i]\right]_{i=l^{*}}^{t^{*}}$, are constructed by sampling $\bm{r}(t)$ and $\bm{u}(t)$ at intervals of length $\tau$ (discrete-time intervals of length 111). $\beta$ is the regularisation parameter and the purpose of the $\beta\,||\textbf{W}_{out}||_{2}^{2}$ term in Eq. (8) is to modify the linear least-squares regression to reduce the magnitudes of elements in $\textbf{W}_{out}$ in order to discourage overfitting.
Minimising Eq. (8) involves computing its partial derivative with respect to $\textbf{W}_{out}$ and setting the resulting expression equal to 0. The $\textbf{W}_{out}$ that minimises Eq. (8) is given by
@@
$\hskip-2.84544pt\dot{\hat{\bm{r}}}(t)$ $=\gamma\left[-\hat{\bm{r}}(t)+\tanh{\left(\textbf{M}\,\hat{\bm{r}}(t)+\sigma\textbf{W}_{in}\textbf{W}_{out}\,\bm{q}(\hat{\bm{r}}(t))\right)}\right],$ (14)
$\hskip-2.84544pt\hat{\bm{r}}\left(0\right)$ $=\bm{r}\left(t_{train}\right),$ (15)
@@ II.3 Training a RC to reconstruct attractors at different parameter settings
$\hskip-5.69046pt\dot{\hat{\bm{r}}}(t\hskip-0.56917pt)\hskip-1.56468pt=\hskip-1.56468pt\gamma\hskip-1.56468pt\left[-\hat{\bm{r}}(t\hskip-0.56917pt)\hskip-0.85355pt+\hskip-0.85355pt\tanh{\left(\textbf{M}\hskip-0.28436pt\hat{\bm{r}}(t\hskip-0.56917pt)\hskip-0.85355pt+\hskip-0.85355pt\sigma\textbf{W}_{in}\textbf{W}_{out}^{C}\,\bm{q}(\hat{\bm{r}}(t\hskip-0.56917pt))\hskip-0.85355pt+\hskip-0.85355pt\bm{b}\right)}\right].$ (20)
@@ II.4 Defining untrained and generated attractors
Definition 1:An attractor, $\pazocal{Z}$, is considered to be an ‘untrained attractor’ if (i) $\pazocal{Z}$ was not present during training, and (ii) $\pazocal{Z}$ does not originate from a continuous deformation/displacement or bifurcation of a given reconstructed attractor by varying certain parameters of the given reservoir computer.
@@
Definition 2:An attractor, $\pazocal{G}$, is considered to be a ‘generated attractor’ if (i) $\pazocal{G}$ was not present during training, and (ii) $\pazocal{G}$ originates from a continuous deformation/displacement or bifurcation of a given reconstructed attractor by varying certain parameters of the given reservoir computer.
@@
We would also like to remark that Definition 2 is designed to compliment Definition 1 in terms of their subtle differences and also to apply to attractors of all types. As a result, Definition 2 may seem relatively trivial in the context of fixed-points and limit cycles. However, we believe that the strength of Definition 2 is its generality.
@@ II.5.1 Task (i) description
$\begin{array}[]{ccl}\dot{x}_{1}&=&10\left(x_{2}+x_{1}\right),\\ \vskip-8.5359pt&&\\ \dot{x}_{2}&=&x_{1}\left(28-x_{3}\right)-x_{2},\\ \vskip-8.5359pt&&\\ \dot{x}_{3}&=&x_{1}x_{2}-\frac{8}{3}x_{3},\end{array}$ (26)
@@
To assess how well the closed-loop RC reconstructs $\pazocal{L}$, we use the method outlined in Appendix B. From this we classify the output from a given initial condition as a ‘good’ reconstruction of $\pazocal{L}$, denoted by $\hat{\pazocal{L}}$, a ‘poor’ reconstruction of $\pazocal{L}$, denoted by $\tilde{\pazocal{L}}$. An example of good and poor reconstructions of $\pazocal{L}$, as determined through the steps outlined in Appendix B, is provided in Fig. 2. If there is no reconstruction of $\pazocal{L}$, i.e. there is no stable attractor located nearby $\pazocal{L}$ with properties resembling $\pazocal{L}$, we then say that the state of the closed-loop RC approaches an UA, corresponding to complete confabulation. We find a wide variation of dynamics corresponding to $\tilde{\pazocal{L}}$ and later show examples of these attractors in Sec. III and Appendix D.
Figure 2: Illustration of decision process in distinguishing between a ‘good’ reconstruction of $\pazocal{L}$ ($\hat{\pazocal{L}}$) in (a) and a ‘bad’ reconstruction of $\pazocal{L}$ ($\tilde{\pazocal{L}}$) in (b). Blue points indicate local maxima that satisfy classification criteria, red point indicate local maxima that do not. Solid black lines indicate a good fit between $\pazocal{L}$ and its reconstruction, dashed black lines indicate a poor fit.
@@
We would like to clarify that the reason why we chose to conduct our experiments with 50 different realisations of M as opposed to 50 different realisations of both M and $\textbf{W}_{in}$ was to simply only vary one component of the RC at a given time. From experiments not presented in the present paper, we found that different UAs can also appear for several different random realisations of $\textbf{W}_{in}$ and a given random M. In future work we intend to investigate the relationship between confabulation dynamics and $\textbf{W}_{in}$ in much greater detail.
@@ II.5.2 Task (ii) description
$\begin{array}[]{ccl}\dot{x}_{1}&=&x_{2}\left(1+x_{3}\right),\\ \vskip-8.5359pt&&\\ \dot{x}_{2}&=&x_{3}\left(x_{2}-ax_{1}\right),\\ \vskip-8.5359pt&&\\ \dot{x}_{3}&=&0.55~x_{3}^{2}-x_{2}^{2}.\end{array}$ (32)
@@
[…]
@@ II.5.3 Task (iii) description
$\begin{array}[]{ccl}\dot{x}_{1}&=&-1.3x_{1}-4\left(x_{2}+x_{3}\right)-x_{2}^{2},\\ \vskip-8.5359pt&&\\ \dot{x}_{2}&=&-1.3x_{2}-4\left(x_{3}+x_{1}\right)-x_{3}^{2},\\ \vskip-8.5359pt&&\\ \dot{x}_{3}&=&-1.3x_{3}-4\left(x_{1}+x_{2}\right)-x_{1}^{2},\end{array}$ (38)
@@ II.6 Summary of related work
While the presence of UAs in RC systems are being increasingly studied and acknowledged [38, 22, 39, 40, 41], and have recently been detected in a physical RC system[42], it is quite likely that UAs have lurked in the background of many researchers dealings with RCs but have simply gone unnoticed. For instance, in one of the most influential papers on reservoir computing in recent times, Lu et al. [23], the authors present an interesting result in their Fig. 5 where they show that for a relatively small change in the input scaling parameter, $\sigma$, there is an abrupt transition in the RC’s dynamics from a period-2 limit cycle to the chaotic Lorenz attractor. Based on the results presented in the present paper, it is not too far-fetched to speculate that the limit cycle is an UA that coexists with the reconstructed Lorenz attractor. Other examples of UAs potentially hiding in plain sight can be seen in de Jong et al. [43].
@@ III.1 Frequency of scenarios in Table 1 across 50 different M
Fig. 5 shows that for $\rho=0$, the only attractor present in each $\mathbb{P}$ is an UA. By increasing $\rho$ to 0.050.050.05, there is a dramatic change in $\mathbb{P}$ as the case of a poor reconstruction of $\pazocal{L}$ coexisting with UAs becomes the dominant scenario. For $\rho=0.1$ we see the first occurrences of RCs that are able to reconstruct $\pazocal{L}$, albeit poorly, without any UAs in $\mathbb{P}$. This trend continues for increasing $\rho$ and when $\rho=0.2$ we see the first instances of RCs that are able to provide a good reconstruction of $\pazocal{L}$ without any UAs present in $\mathbb{P}$. By increasing $\rho$ further we see this becomes the dominant scenario for $0.3\lesssim\rho\lesssim 0.7$. Within this range of $\rho$ values, UAs are found for a decreasing number of M. For $0.7\lesssim\rho\lesssim 1.1$, Fig. 5 shows that there is a relatively equal likelihood that the RC will provide either a good or bad reconstruction of $\pazocal{L}$, and that it is much less likely for these reconstructions to coexist with UAs. However, there is a steady increase in the number of RCs that fail to provide any reconstruction of $\pazocal{L}$ within this range of $\rho$ values, thus, resulting in an increase in the UA only scenario. These trends continue for $\rho>1.1$ with poor reconstructions of $\pazocal{L}$ becoming the dominant scenario.
@@
In Fig. 8 we re-present the results of Fig. 5 from the point of view of each M in order to illustrate the changes that occur in each of the corresponding $\mathbb{P}$ when changing $\rho$. We label each M by $\textbf{M}_{i}$, from $i=1$ to 505050, in order of the number of $\rho$ values that the corresponding closed-loop RC in Eq. (14) provides a good reconstruction of $\pazocal{L}$, with $i=1$ being the most amount of $\rho$ values. Fig. 8 provides us with additional information, such as: some random realisations of M quickly overcome the influence of UAs, like $\textbf{M}_{1}$ and $\textbf{M}_{4}$, whereas others require relatively large $\rho$, like $\textbf{M}_{43}$ needing $\rho>0.9$.
@@ III.2 UAs filling in the gaps when reconstruction of $\pazocal{L}$ fails
Fig. 8 shows that when $\rho=0.05$ the majority of the different M matrices are able to provide a poor reconstruction of $\pazocal{L}$, denoted by $\tilde{\pazocal{L}}$, which coexists with at least one UA. In this subsection, we illustrate an interesting route in which the RC reconstructs $\pazocal{L}$ that involves period-doubling bifurcations. In Appendix C we illustrate the most common way that the RC reconstructs $\pazocal{L}$ when increasing $\rho$ from 0. In Appendix D we provide additional examples of interesting ways in which the RC reconstructs $\pazocal{L}$.
III.2.1 Reconstruction of $\pazocal{L}$ via period-doubling route to chaos